The sum of two numbers is one hundred seventy eight. The difference between the number is six. What is the product of the two numbers ?
Answer by stanbon(75887)

The sum of two numbers is one hundred seventy eight. The difference between the number is six. What is the product of the two numbers ?
--------
Equations:
x + y = 78
x - y = 6
------------------
Add to get::
2x = 84
x = 42
----
Solve for "y":
x + y = 78
42 + y = 78
y = 36
=================
Cheers,
Stan H.